THE SECOND BOUNDED COHOMOLOGY OF

A GROUP WITH INFINITELY MANY ENDS

Koji Fujiwara

Abstract. We study the second bounded cohomology of an amalgamated free product of groups, and an HNN extension of a group. As an application, we have a group with infinitely many ends has infinite dimensional second bounded cohomology.

Typeset by AMS-TEX